OpenStudy (dayianaforever):

Solve the system by elimination. -2x+2y+3z=0 -2x-y+z=-3 2x+3y+3z=5 i just need to know how to do it i honestly can do it by my self need help!!!!

OpenStudy (dayianaforever):

please anyone

rishavraj (rishavraj):

see add up eqation 1 and 3 .....x gets eliminated.... and subtract equation 2 from 1 ... x gets eliminated ...

rishavraj (rishavraj):

http://prntscr.com/7om5on

OpenStudy (anonymous):

The elimination method for solving systems of linear equations uses the addition property of equality. You can add the same value to each side of an equation. So if you have a system: x – 6 = −6 and x + y = 8, you can add x + y to the left side of the first equation and add 8 to the right side of the equation. And since x + y = 8, you are adding the same value to each side of the first equation.

rishavraj (rishavraj):

or else multiply equation 2 from -1 and then add it with equation 1....x gets eliminated

OpenStudy (dayianaforever):

ohh okay thank you @rishavraj and @NoahDabreo